# EDUMATH: Generating Standards-aligned Educational Math Word Problems

### v1

@@ Abstract
Math word problems (MWPs) are critical K-12 educational tools, and customizing them to students’ interests and ability levels can increase learning outcomes. However, teachers struggle to find time to customize MWPs for each student given large class sizes and increasing burnout. We propose that LLMs can support math education by generating MWPs customized to student interests and math education standards. To this end, we use a joint human expert-LLM judge approach to evaluate over 11,000 MWPs generated by open and closed LLMs and develop the first teacher-annotated dataset for standards-aligned educational MWP generation. We show the value of our data by using it to train a 12B open model that matches the performance of larger and more capable open models. We also use our teacher-annotated data to train a text classifier that enables a 30B open LLM to outperform existing closed baselines without any training. Next, we show our models’ MWPs are more similar to human-written MWPs than those from existing models. We conclude by conducting the first study of customized LLM-generated MWPs with grade school students, finding they perform similarly on our models’ MWPs relative to human-written MWPs but consistently prefer our customized MWPs.  [^1] Our code, models, and data can be found at https://github.com/bryanchrist/EDUMATH
@@
University of Virginia brc4cb@virginia.edu ,fnt7et@virginia.edu, jk8sd@virginia.edu, hartvigsen@virginia.edu
@@ 1 Introduction
Math word problems (MWPs) are natural language math questions paired with numerical solutions, and are important elements of K-12 math education because they are an effective tool to verify mastery of a specific math concept Daroczy et al. (2015); Pearce et al. (2013); Schwartz (2023); Verschaffel et al. (2020). In all K-12 grades, personalizing MWPs to students’ interests and ability levels can support learning outcomes by increasing their interest in math and ensuring problems are within their readiness level Baker et al. (2020); Bernacki and Walkington (2018); Nasir and Hand (2008); Pinkard et al. (2020); Walkington (2013). Typically, teachers manually write or curate MWPs customized to students’ interests or ability levels. However, with competing responsibilities, limited time, and increasing burnout Kuncl and Christensen (2024), teachers must often rely on standard question sets from educational websites or practice tests for end-of-year examinations, which are limited in quantity, not customized, and not always easily searchable for math standards/topics VDOE . We propose that large language models (LLMs) can address this challenge by automatically generating MWPs aligned with students’ interests and ability levels given their vast natural language ability and increasing abilities in math reasoning Ahn et al. (2024).
However, given recent work has found LLMs struggle to generate educational MWPs Ariyarathne et al. (2025); Christ et al. (2024), it is important to conduct further study before directly using them in educational settings. To this end, we aim to explore and enhance LLMs’ capacity to generate educational MWPs tailored to students’ interests and ability levels as indicated by alignment with math standards from the United States. Recent work has explored generating educational MWPs with LLMs customized to students’ interests Christ et al. (2024), but not math standards. Other work has explored generating math practice problems, including MWPs, with LLMs that are aligned with math topics Ariyarathne et al. (2025); Sun et al. (2025), but they only generate problems without associated solutions. Two other works explore generating MWPs with LLMs using reference problems Niyarepola et al. (2022); Zong and Krishnamachari (2023), but they too only generate problems without solutions. To be useful to students and teachers, generated problems must have associated solutions written in clear natural language and be aligned with particular math standards (e.g., single-step whole number multiplication not exceeding 100) rather than general math topics like multiplication to allow teachers fine-grained control over generated problems. We refer to this task as standards-aligned educational MWP generation.
@@
We conduct this study in five phases. In phase one, we use a joint expert-LLM judge approach to annotate a small existing human-written MWP dataset for grade school math standards given there are no datasets fully annotated for these standards. In phase two, we use the labeled dataset to prompt an existing LLM to generate over 3,000 MWPs that teachers annotate for our four criteria to expand the amount of training and prompting data available for this task. In phase three, we train a mid-size (12B) LLM using annotation data from the second phase to show the value of our data for training standards-aligned educational MWP generators. In this stage, we further show the quality of our annotated data by using it to train a simple text classifier that filters outputs from both our trained model and an existing 30B LLM without further training. In phase four, we evaluate over 8,000 MWPs from our 12B and 30B models as well as several open and closed baselines with an LLM judge aligned with feedback from teachers. In the final phase, we conduct the first evaluation of customized LLM-generated MWPs with real grade school students.
Through these evaluations, we find a performance gap in the task of generating standards-aligned educational MWPs between open and closed models and between smaller and larger open models. We show the value of our data by demonstrating both that training on it is sufficient to nearly close the gap in performance between smaller and larger open LLMs and filtering outputs using our trained text classifier is enough to help a 30B LLM outperform closed baselines without any training. We also show that our data and models’ outputs are higher quality and more similar to existing human-written data than outputs from comparison models on a wide range of automatic metrics. Finally, we show that students prefer our generated MWPs to those that are human-written while performing similarly on them. Our key contributions are:
  • We find existing open models, especially smaller ones, struggle at the task of generating standards-aligned educational MWPs relative to closed baselines.
@@ 2 Related Work
There are several popular datasets for training and evaluating LLMs on grade school math, including GSM8K Cobbe et al. (2021), NumGLUE Mishra et al. (2022), GSM-Hard Gao et al. (2023), ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021). Many works have also developed additional grade school math benchmark datasets or large training datasets consisting of different solution rationales for existing datasets, particularly GSM8K, including both those that are human written Kim et al. (2023); Mishra et al. (2023) and synthetic Mitra et al. (2024); Shi et al. (2023); Toshniwal et al. (2024); Yu et al. (2023); Yuan et al. (2023); Yue et al. (2023).
Because existing datasets are designed to train and evaluate grade school math reasoning, they are not aligned with training standards-aligned educational MWP generators. Training data for standards-aligned educational MWP generators must contain high-quality grammar, have both questions and answers, contain readable solutions for teachers and students, be segmented by grade level and standards, have questions that are appropriate for the given grade levels and standards, have questions that are similar to those that are written by educators, and contain content appropriate for a school setting. We consider high-quality grammar and similarity to MWPs students would see in a classroom setting as minimally sufficient criteria any potentially useful training dataset must possess. Existing datasets that contain these minimum sufficiency criteria and are aligned with one or more of the other criteria defined above are GSM-Hard Gao et al. (2023), GSM8K Cobbe et al. (2021), MathInstruct GSM8K Yue et al. (2023), ASDIV Miao et al. (2020), SVAMP Patel et al. (2021), EGSM Christ et al. (2024), and Mathwizards Ariyarathne et al. (2025), although none of these datasets contain all criteria.
@@
Closest to our work is Ariyarathne et al. (2025), who use LLMs to generate elementary school MWPs aligned with a particular math topic like single-digit addition. However, they do not automatically generate solutions, nor do they align their problems with the full language of math standards, which specify both topics and complexity constraints MWPs must follow (e.g., single-step single-digit addition). In addition, they do not conduct evaluations involving real teachers and, instead, rely on automated evaluations and evaluations from non-education college students. Another closely related work is Sun et al. (2025), who use LLMs to generate math problems aligned with a math topic and difficulty level. However, they do not automatically generate solutions, nor are their problems aligned with math standards; instead, they are aligned with general math topics like the Pythagorean theorem. Further, they do not involve teachers in their evaluations and instead rely only on automatic evaluation methods.
Most other MWP generation works use LLMs or traditional NLP approaches to generate MWPs based on pre-specified equations, re-write existing MWPs, or provide additional MWPs based on reference problems Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Kumar and Sreenivasa Kumar (2025); Norberg et al. (2023); Niyarepola et al. (2022); Qin et al. (2023, 2024); Wang et al. (2021); Wu et al. (2022); Zhou and Huang (2019); Zong and Krishnamachari (2023); Zhou et al. (2023), each of which limit the originality and range of outputs and can often lead to rephrasing training data. These approaches also only generate problems without solutions and require additional input from users, which is infeasible for teachers or students who want to generate customized MWPs, making them incomparable to our work. To address these issues, we generate MWP question/answer pairs based on a grade level and standard simultaneously without requiring input reference problems or equations.
@@ 3 Methods
We define math word problems (MWPs) as natural language questions describing contextual math scenarios paired with solutions written in a natural language, chain-of-thought (CoT) Wei et al. (2023) format. We study generating MWPs with LLMs aligned with a math standard provided in the prompt, as shown in Figure 1. We evaluate generated MWPs based on the grade levels, standards, evaluation criteria, and automatic metrics defined below and conduct a multi-stage pipeline to generate training data and train models for the task of standards-aligned educational MWP generation.
@@
We study generating MWPs for grades 3-5 in the US. We choose to study grades 3-5 for two main reasons: 1) choosing a narrow grade range allows us to fully assess models’ abilities to generate MWPs for all relevant math standards in these grade levels, as each grade has many standards that can be assessed with MWPs, and 2) these are the grade levels for which MWPs are most used and, therefore, where a standards-aligned educational MWP generator would be most useful Daroczy et al. (2015); Pearce et al. (2013); Schwartz (2023); Verschaffel et al. (2020). We use MWP math standards from the Virginia Standards of Learning (VA SOL) because they reference bounded difficulty levels and number ranges for all standards, while the Common Core standards, which are used in many US states, are not always as specific Initiative ; Virginia Department of Education . We provide a mapping of VA SOLs to Common Core standards in the repo associated with this paper.
@@
We evaluate MWPs on four criteria. The first three criteria, Solvability, Accuracy, and Educational Appropriateness, are adapted from Christ et al. (2024). Solvability evaluates if questions are possible to solve and have one correct answer. Accuracy assesses if generated CoT solutions are accurate. We label MWPs with accurate final answers but inaccurate, confusing, or overly complicated reasoning as inaccurate because our solutions must be readable for teachers and students. Educational Appropriateness evaluates if teachers would feel comfortable giving the MWP to one of their students. In general, an educationally appropriate MWP is one that makes sense, does not contain grammar errors or conflicting information, and is about topics that are appropriate for 3rd-5th grade students in a school setting. The final criterion we introduce in this work is Standards Alignment, or if the MWP effectively incorporates a pre-specified math standard. Because the latter three criteria are hard to precisely define, we follow Christ et al. (2024) in emphasizing teacher involvement in evaluating MWPs. See Appendix G for MWPs that do and do not meet these criteria.
@@
Readability automatically assesses the appropriateness of MWPs. We use a composite of 8 readability metrics as implemented by Georgia State University’s ARTE tool GSU to comprehensively assess readability: the Automated Readability Index (ARI) Kincaid et al. (1975), Flesch-Kincaid Grade Level (FKGL) Flesch (1948), Flesch Reading Ease (FRE) Kincaid et al. (1975), CAREC and CAREC-M Crossley et al. (2019), Coh-Metrix L2 Reading Index (CML2RI) Crossley et al. (2008), New Dale-Chall Chall and Dale (1995), and the Simple Measure of Gobbledygook (SMOG) Mc Laughlin (1969). We use min-max scaling to normalize each metric in the range of [0, 1] and flip the CML2RI and FRE scores so that higher scores represent more difficult to read texts to match the other metrics. We then report the average of these scaled metrics to capture the overall reading complexity of each MWP. In general, lower reading levels are preferred for MWPs because high reading levels can harm student performance, especially for those struggling in math Walkington et al. (2018). We also calculate each MWP’s average token length, which is a proxy for complexity given that longer questions often include more mathematical operations.
Following prior work Christ et al. (2024); Jiao et al. (2023); Zhou et al. (2023), we use Perplexity (PPL) and BERTScore to evaluate the quality of our synthetic MWPs and compare them to human-written MWPs, respectively. Lower PPL suggests higher-quality outputs and we calculate PPL with GPT-2. We use BERTScore Zhang et al. (2020) to compute the semantic similarity of generated MWPs to each other and compare it to that of human-written MWPs. A higher BERTScore for generated MWPs suggests they are more similar to each other than human-written MWPs, and a lower BERTScore suggests the opposite. We also compute BERTScore for generated MWPs relative to human-written MWPs to identify if they are similar and use the BERTScore of human-written MWPs relative to each other as a reference to compare against. For all comparisons to human-written MWPs, we use the 3rd-5th grade subset of ASDIV since it is the only human-written English MWP dataset labeled for grade level.
@@ 3.1 Labeling Existing Data for Math Standards
MWPs with labels for math standards are required for training or prompting models to generate standards-aligned educational MWPs. However, no grade school math question/answer datasets are labeled for standards. ASDIV Miao et al. (2020) has labels for grade and math topics like addition, but not standards, so we used a four-stage approach to label its 3rd-5th grade MWPs (n = 1,027) for VA SOLs. In stage one, an undergraduate education student labeled MWPs for all relevant VA SOLs, as MWPs may address more than one standard. In stage two, a research team member with K-12 teaching experience checked the labels and made adjustments, often reducing the number of VA SOLs associated with MWPs in cases where the student over-labeled MWPs. We also identified and removed two MWPs that were unsolvable and rewrote several others to make them solvable, as some referenced constraints described in previous MWPs that are not included in the current MWP.
In stage three, we prompted Llama 3.3 70B IT Meta with each ASDIV MWP and its VA SOLs from stage two and asked if the question included each VA SOL. See Appendix I.1 for the prompt. We also used a few-shot prompt (see Appendix I.2) for generating readable CoT solutions for the ASDIV subset given it does not have solutions in this format. We compared generated solutions against ground-truth answers to ensure solution accuracy. In stage four, our research team member with K-12 teaching experience reviewed the Llama 3.3 standard labels and made adjustments to remove VA SOLs that were not incorporated in MWPs. To ensure solution quality, we had Gemma 3 27B IT Team et al. (2025) review generated solutions and rewrite any it deemed were too complex or contained incorrect/unnecessary intermediate reasoning (see Appendix I.3 for the prompt). Finally, the undergraduate education student reviewed solutions to ensure intermediate reasoning steps were accurate and the solution was readable for teachers/students. This resulted in a final dataset with 1,025 MWPs with gold labels for VA SOLs and CoT solutions (see Appendix H for an example).

### v2

@@ Abstract
Math word problems (MWPs) are critical K-12 educational tools, and customizing them to students’ interests and ability levels can enhance learning. However, teachers struggle to find time to customize MWPs for students given large class sizes and increasing burnout. We propose that LLMs can support math education by generating MWPs customized to student interests and math education standards. We use a joint human expert-LLM judge approach to evaluate over 11,000 MWPs generated by open and closed LLMs and develop the first teacher-annotated dataset for standards-aligned educational MWP generation. We show the value of our data by using it to train a 12B open model that matches the performance of larger and more capable open models. We also use our teacher-annotated data to train a text classifier that enables a 30B open LLM to outperform existing closed baselines without any training. Next, we show our models’ MWPs are more similar to human-written MWPs than those from existing models. We conclude by conducting the first study of customized LLM-generated MWPs with grade school students, finding they perform similarly on our models’ MWPs relative to human-written MWPs but consistently prefer our customized MWPs.  [^1] Our code, models, and data can be found at https://github.com/bryanchrist/EDUMATH
@@
University of Virginia Correspondence: brc4cb@virginia.edu
@@ 1 Introduction
Math word problems (MWPs) are natural language math questions paired with numerical solutions, and are critical elements of K-12 math education because they help verify mastery of a math concept Daroczy et al. (2015); Pearce et al. (2013); Schwartz (2023); Verschaffel et al. (2020). In all K-12 grades, personalizing MWPs to students’ interests and ability levels supports learning by increasing their interest in math and ensuring problems are within their readiness level Baker et al. (2020); Bernacki and Walkington (2018); Nasir and Hand (2008); Pinkard et al. (2020); Walkington (2013). Typically, teachers manually write or curate MWPs customized to students’ interests or ability levels. However, with competing responsibilities, limited time, and increasing burnout Kuncl and Christensen (2024), teachers must often rely on standard question sets from educational websites or practice tests for end-of-year examinations, which are limited in quantity, not customized, and not always easily searchable for math standards/topics VDOE . We propose that large language models (LLMs) can address this challenge by automatically generating MWPs aligned with students’ interests and ability levels given their vast natural language ability and increasing level of math reasoning Ahn et al. (2024).
However, given recent work has found LLMs struggle to generate educational MWPs Ariyarathne et al. (2025); Christ et al. (2024), further study must be conducted before directly using them in educational settings. We aim to explore and enhance LLMs’ capacity to generate educational MWPs tailored to students’ interests and ability levels as indicated by alignment with math standards from the US. Recent work has generated educational MWPs with LLMs customized to students’ interests Christ et al. (2024), but not math standards. Other work has generated math practice problems, including MWPs, with LLMs aligned with math topics Ariyarathne et al. (2025); Sun et al. (2025), but they only generate problems without solutions. Two other works have generated MWPs with LLMs using reference problems Niyarepola et al. (2022); Zong and Krishnamachari (2023), but they too generate problems without solutions. To be useful to students and teachers, generated problems must have solutions written in clear natural language and be aligned with a math standard (e.g., single-step whole number multiplication not exceeding 100) rather than general math topics like multiplication to allow teachers fine-grained control over output MWPs. We refer to this task as standards-aligned educational MWP generation.
@@
Our study has five phases. In phase one, we use a joint expert-LLM judge approach to annotate a small human-written MWP dataset for grade school math standards given there are no datasets fully annotated for these standards. In phase two, we use the labeled data to prompt a LLM to generate over 3,000 MWPs that teachers annotate for our four criteria to expand the amount of training and prompting data for this task. In phase three, we train a mid-size (12B) LLM using the second phase’s annotation data to validate our data for training standards-aligned educational MWP generators. In this stage, we further show the quality of our annotated data by using it to train a small text classifier that filters outputs from both our trained model and an existing 30B LLM without further training. In phase four, we evaluate over 8,000 MWPs from our models as well as several open and closed baselines with an LLM judge aligned with feedback from teachers. In phase five, we conduct the first study of customized LLM-generated MWPs with real grade school students.
Through our evaluations, we find performance gaps in the task of generating standards-aligned educational MWPs between open and closed models and between smaller and larger open models. We validate our data by demonstrating both that training on it is sufficient to nearly close the gap in performance between smaller and larger open LLMs and filtering outputs using our text classifier is enough to help a 30B LLM outperform closed baselines without any training. We also show our data and models’ outputs are higher quality and more similar to human-written MWPs than outputs from comparison models on several automatic metrics. Finally, we show students prefer our generated MWPs to human-written MWPs while performing similarly on them. Our key contributions are:
  • We find open models, especially small ones, struggle to generate standards-aligned educational MWPs relative to closed models.
@@ 2 Related Work
There are several popular datasets for training and evaluating LLMs on grade school math, including GSM8K Cobbe et al. (2021), GSM-Hard Gao et al. (2023), ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021). Many works have developed other grade school math benchmark datasets or large training datasets consisting of different solution rationales for existing datasets, particularly GSM8K, including both those that are human written Kim et al. (2023); Mishra et al. (2023) and synthetic Li et al. (2024a, b); Mitra et al. (2024); Shi et al. (2023); Tang et al. (2024); Toshniwal et al. (2024); Yu et al. (2023); Yuan et al. (2023); Yue et al. (2023).
Because existing datasets are designed to train and evaluate grade school math reasoning, they are not aligned with training standards-aligned educational MWP generators. Training data for such generators must contain high-quality grammar, have both questions and answers, contain readable solutions for teachers and students, be segmented by grade level and standards, have questions that are appropriate for the given grade levels and standards, have questions similar to teacher-written questions, and contain content appropriate for a school setting. We consider high-quality grammar and similarity to MWPs students see in a classroom setting as minimally sufficient criteria any potentially useful training dataset must possess. Existing datasets that contain these minimum sufficiency criteria and are aligned with one or more of the other criteria defined above are GSM-Hard Gao et al. (2023), GSM8K Cobbe et al. (2021), MathInstruct GSM8K Yue et al. (2023), ASDIV Miao et al. (2020), SVAMP Patel et al. (2021), EGSM Christ et al. (2024), and Mathwizards Ariyarathne et al. (2025), although none contain all criteria.
@@
Closest to our work is Ariyarathne et al. (2025), who use LLMs to generate elementary MWPs aligned with a specific math topic like single-digit addition. However, they do not generate solutions, nor do they align their problems with the full language of math standards, which specify both topics and complexity constraints MWPs must follow (e.g., single-step single-digit addition). Further, they do not conduct evaluations involving real teachers and, instead, rely on automated evaluations and evaluations with non-education college students. Another closely related work is Sun et al. (2025), who use LLMs to generate math problems aligned with a math topic and difficulty level. However, they do not generate solutions, nor are their problems aligned with math standards; instead, they are aligned with general math topics like the Pythagorean theorem. Further, they do not involve teachers in their evaluations and instead rely only on automatic evaluation methods.
Most other MWP generation works use LLMs or traditional NLP approaches to generate MWPs based on pre-specified equations, re-write existing MWPs, or provide new MWPs based on reference problems Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Kumar and Sreenivasa Kumar (2025); Norberg et al. (2023); Niyarepola et al. (2022); Qin et al. (2023, 2024); Wang et al. (2021); Wu et al. (2022); Zhou and Huang (2019); Zong and Krishnamachari (2023); Zhou et al. (2023), each of which limit the originality and range of outputs and can often lead to rephrasing training data. These approaches also only generate problems without solutions and require additional input from users, which is infeasible for teachers or students who want to generate customized MWPs, making them incomparable to our work. To address these issues, we generate MWP question/answer pairs based on a grade level and standard simultaneously without requiring input reference problems or equations.
@@ 3 Methods
We define math word problems (MWPs) as natural language questions describing contextual math scenarios paired with solutions written in a natural language, chain-of-thought (CoT) Wei et al. (2023) format. We study generating MWPs with LLMs aligned with a math standard provided in the prompt, as shown in Figure 1. We evaluate generated MWPs based on the grade levels, standards, evaluation criteria, and automatic metrics defined below and conduct a multi-stage pipeline to generate training data and train models for the task of standards-aligned educational MWP generation (see Appendix A for a visual of this pipeline).
@@
We study generating MWPs for grades 3-5 in the US. We choose to study grades 3-5 for two reasons: 1) choosing a narrow grade range allows us to fully assess models’ abilities to generate MWPs for all relevant math standards in these grade levels, as each grade has many standards that can be assessed with MWPs, and 2) these are the grade levels for which MWPs are most used and, therefore, where a MWP generator would be most useful Daroczy et al. (2015); Pearce et al. (2013); Schwartz (2023); Verschaffel et al. (2020). Expanding to other grade levels is a promising direction for future work but would dilute teacher annotation efforts given a fixed annotation budget. We use MWP math standards from the Virginia Standards of Learning (VA SOL) because they reference bounded difficulty levels and number ranges for all standards, while the Common Core standards, which are used in many US states, are not always as specific Initiative ; Virginia Department of Education . For all grades, VA SOLs are hierarchical and build on one another such that standards encountered later rely on those learned previously Virginia Department of Education . We provide a mapping of VA SOLs to Common Core standards in this paper’s repo, enabling future work to switch between these standard sets without needing to retrain our models.
@@
We evaluate MWPs on four criteria. The first three criteria are adapted from Christ et al. (2024). Solvability evaluates if questions are possible to solve and have one correct answer. Accuracy assesses if generated CoT solutions are accurate. We label MWPs with accurate final answers but inaccurate, confusing, or overly complicated reasoning as inaccurate because our solutions must be readable for teachers and students. Educational Appropriateness evaluates if teachers be comfortable giving the MWP to one of their students. An educationally appropriate MWP is one that makes sense, does not contain grammar errors or conflicting information, and is about topics appropriate for 3rd-5th grade students in a school setting. The final criterion we introduce in this work is Standards Alignment, or if the MWP effectively incorporates a pre-specified math standard. While the latter three criteria are hard to precisely define, they are similar to those used in other MWP generator works Christ et al. (2024); Jiao et al. (2023); Qin et al. (2023, 2024) and are meant to model how real teachers evaluate MWPs, which is why emphasize teacher involvement in evaluating MWPs. See Appendix H for MWPs that do and do not meet these criteria.
@@
Readability automatically assesses the appropriateness of MWPs. We use a composite of 8 readability metrics to comprehensively assess readability: the Automated Readability Index (ARI) Kincaid et al. (1975), Flesch-Kincaid Grade Level (FKGL) Flesch (1948), Flesch Reading Ease (FRE) Kincaid et al. (1975), CAREC and CAREC-M Crossley et al. (2019), Coh-Metrix L2 Reading Index (CML2RI) Crossley et al. (2008), New Dale-Chall Chall and Dale (1995), and the Simple Measure of Gobbledygook (SMOG) Mc Laughlin (1969). We use min-max scaling to normalize each metric in the range of [0, 1] and flip the CML2RI and FRE scores so that higher scores represent more difficult to read texts to match the other metrics. We then report the average of these scaled metrics to capture the overall reading complexity of each MWP. In general, lower reading levels are preferred for MWPs because high reading levels can harm student performance, especially for those struggling in math Walkington et al. (2018). We also calculate each MWP’s average token length, which is a proxy for complexity given that longer questions often include more mathematical operations.
Like prior work Christ et al. (2024); Jiao et al. (2023); Zhou et al. (2023), we use Perplexity (PPL) and BERTScore to evaluate the quality of our synthetic MWPs and compare them to human-written MWPs, respectively. Lower PPL suggests higher-quality MWPs, which we calculate with GPT-2. We use BERTScore Zhang et al. (2020) to compute the semantic similarity of synthetic MWPs to each other and compare it to that of human-written MWPs. A higher BERTScore for synthetic MWPs suggests they are more similar to each other than human-written MWPs, and a lower BERTScore suggests the opposite. We compute BERTScore for synthetic MWPs relative to human-written MWPs to identify if they are similar and use the BERTScore of human-written MWPs relative to each other as a reference. For all comparisons to human-written MWPs, we use ASDIV’s 3rd-5th grade subset since it is the only human-written English MWP dataset labeled for grade level.
@@ 3.1 Labeling Existing Data for Math Standards
MWPs with labels for math standards are required for training or prompting models to generate standards-aligned educational MWPs. However, no grade school math question/answer datasets are labeled for standards. ASDIV Miao et al. (2020) has labels for grade and math topics like addition, but not standards, so we used a four-stage approach to label its 3rd-5th grade MWPs (n = 1,027) for VA SOLs. In stage one, an undergraduate education student labeled MWPs for all relevant VA SOLs, as MWPs may address more than one standard. In stage two, a research team member with K-12 teaching experience checked the labels and made adjustments, often reducing the number of VA SOLs associated with MWPs in cases where the student over-labeled MWPs. We also identified and removed two MWPs that were unsolvable and rewrote several others to make them solvable.
In stage three, we prompted Llama 3.3 70B IT Meta with each ASDIV MWP and its VA SOLs from stage two and asked if the question included each VA SOL. See Appendix J.1 for the prompt. We also used a few-shot prompt (see Appendix J.2) for generating readable CoT solutions for the ASDIV subset given it does not have solutions in this format. We compared generated solutions against ground-truth answers to ensure solution accuracy. In stage four, our research team member with K-12 teaching experience reviewed the Llama 3.3 standard labels and made adjustments to remove VA SOLs that were not incorporated in MWPs. To ensure solution quality, we had Gemma 3 27B IT Team et al. (2025) review generated solutions and rewrite any it deemed were too complex or contained incorrect/unnecessary intermediate reasoning (see Appendix J.3 for the prompt). Finally, the undergraduate education student reviewed solutions to ensure intermediate reasoning steps were accurate and the solution was readable for teachers/students. This resulted in a final dataset with 1,025 MWPs with gold labels for VA SOLs and CoT solutions (see Appendix I for an example).